# Not All Heads Matter: A Head-Level KV Cache Compression Method with Integrated Retrieval and Reasoning

### v3

@@ Abstract
Key-Value (KV) caching is a common technique to enhance the computational efficiency of Large Language Models (LLMs), but its memory overhead grows rapidly with input length. Prior work has shown that not all tokens are equally important for text generation, proposing layer-level KV cache compression to selectively retain key information. Recognizing the distinct roles of attention heads in generation, we propose HeadKV, a head-level KV cache compression method, and HeadKV-R2, which leverages a novel contextual reasoning ability estimation for compression. Our approach operates at the level of individual heads, estimating their importance for contextual QA tasks that require both retrieval and reasoning capabilities. Extensive experiments across diverse benchmarks (LongBench, LooGLE), model architectures (e.g., Llama-3-8B-Instruct, Mistral-7B-Instruct), and long-context abilities tests demonstrate that our head-level KV cache compression significantly outperforms strong baselines, particularly in low-resource settings (KV size = 64 & 128). Notably, our method retains just 1.5% of the KV cache while achieving 97% of the performance of the full KV cache on the contextual question answering benchmark.  [^1] https://github.com/FYYFU/HeadKV/
@@ 1 Introduction
We conduct experiments on various benchmarks requiring both retrieval and reasoning abilities, including QA tasks from LongBench (Bai et al., 2024) and LooGLE (Li et al., 2024b), using backbone models such as Llama-3-8B-Instruct (Dubey et al., 2024) and Mistral-7B-Instruct (Jiang et al., 2023). Our results demonstrate that the head-level KV cache compression method outperforms previous approaches on nearly all tasks. Furthermore, the allocation strategy based on our estimated importance scores—reflecting both retrieval and reasoning abilities—outperforms allocation strategies based on retrieval ability alone. In challenging scenarios like needle-in-a-Haystack and reasoning-in-a-Haystack tests, our methods effectively preserve the model’s retrieval and reasoning capabilities. Finally, experiments on memory and latency reveal that our approach significantly reduces memory usage and decoding latency, comparing with the original full KV.
@@ 2.2 KV cache compression
Improving computational efficiency for LLMs, particularly for extremely long inputs, has attracted considerable research interest (Shazeer, 2019; Chen et al., 2023; Dao, 2024; Zhou et al., 2024; Ye et al., 2024), including caching key and value vectors in Multi-Head Attention to improve generation efficiency (Xiao et al., 2024; Zhang et al., 2023; Li et al., 2024d; Cai et al., 2024). One challenge with full KV cache approaches is that the memory usage increases linearly with input length, leading to significant memory management challenges. Various KV cache compression techniques have been proposed to reduce memory usage and improve inference efficiency. For example, Xiao et al. (2024) addressed the ‘attention sink’ issue with StreamingLLM, retaining only the first $k$ tokens’ KV cache. Zhang et al. (2023) applied the Heavy Hitter Oracle strategy to select key cache entries, while Li et al. (2024d) used the attention scores of the last $\alpha$ tokens to identify relevant entries. Cai et al. (2024) introduced PyramidKV, assigning smaller cache budgets to higher layers based on attention matrix patterns.
@@ 3.1 Head-level Importance Score Estimation
$S_{h}=\sum_{t=1}^{N}\mathcal{N}^{t},\quad\text{where}\ \mathcal{N}^{t}=\begin{cases}\frac{1}{N}&\text{if }\arg max({\bm{a}}_{h})\in{\bm{k}}\\ 0&\text{otherwise}\\ \end{cases}$ (1)
where $N$ is the length of the inserted answer ${\bm{k}}$, and ${\bm{a}}_{h}$ is the attention score on the combined input from head $h$. By using various settings of the inserted position $p_{i}$, they obtain the head-level Retrieval Heads distribution.
@@
$S_{h}=\sum_{t=1}^{N}\sum_{i=1}^{N}\mathcal{M}^{t}_{i},\quad\text{where}\ \mathcal{M}^{t}_{i}=\begin{cases}\frac{{\bm{a}}_{i}}{N}&\text{if }\text{top-}i({\bm{a}}_{h})\in{\bm{c}}^{2}\\ 0&\text{otherwise}\\ \end{cases}$ (2)
@@ 3.2 Head-Level KV Cache Allocation
Building on the head-level importance distributions, we propose a comprehensive KV cache allocation strategy. Each head $h$ is initially assigned a fixed KV cache size $b$ with an associated importance score $S_{h}$. To allow dynamic allocation, we create a shared budget pool $B$ by extracting a portion of the budget from each head, leaving the remainder as the basic budget. This process is illustrated in the Head-Level Allocation section of Figure 1. The budget pool $B$ is then distributed among the heads in proportion to their importance scores $S_{h}$. The final head-level KV cache allocation is as follows:
@@ 4.1 Experiment settings
We compare our head-level KV cache compression method against strong baselines using two open-source LLMs: Llama-3-8B-Instruct (Dubey et al., 2024) and Mistral-7B-Instruct (Jiang et al., 2023). The evaluation is based on two benchmarks for long-context understanding: LongBench (Bai et al., 2024) and LooGLE (Li et al., 2024b). For LongBench, we use datasets from the Single-Doc QA and Multi-Doc QA categories to assess contextual reasoning. For LooGLE, we focus on the Long Dependency QA task, which includes four QA-related tasks. Dataset details are in Appendix Table 5.
@@ 4.4 Long-Context Retrieval and Reasoning
Figure 5: Neele-in-a-Haystack test results on Llama-3-8B-Instruct with KV cache = 128. We build our head-level KV cache method based on SnapKV and our proposed method significantly outperform all strong baselines. Moreover, our Retrieval-Reasoning Heads distribution maintains and improves long context retrieval ability. Results on Mistral-7B-Instruct can be found in Appendix Figure 8, which are consistent with results on Llama-3-8B-Instruct.
@@ 4.5 Memory & Latency
To assess the decoding latency of each method, we use 32K-length data from the Reasoning-in-a-Haystack test as input and set various generation lengths (1, 512, 1024, 2048, 4096) for comparison. As shown in the Decoding Latency of Figure 6, our proposed method achieves the same decoding latency as other KV cache compression methods while maintaining performance closest to the Full KV cache. Notably, the decoding time includes both the pre-filling time and the decoding time. Therefore, we can conclude that the pre-filling time for our method and other baselines (such as PyramidKV, Ada-SnapKV) is almost negligible, as shown at the starting point (generation length = 1) of Decoding Latency in Figure 6.
@@ Appendix A Detaile Results
Method Single-Doc QA Multi-Doc QA Avg. Long dependency QA Avg.
@@
Llama-3-8B-Instruct, KV Size=Full
FKV 25.56 32.07 39.71 43.57 35.28 21.18 32.90 8.73 11.21 0.67 7.43 7.01
@@
SKV 20.51 12.80 31.69 37.02 25.91 17.02 24.16 8.84 9.43 0.66 6.18 6.28
PyramidKV 21.17 13.66 29.34 34.86 23.46 15.88 23.06 8.27 9.31 0.63 6.86 6.27
Ada-SKV 22.26 17.30 33.37 39.82 27.86 17.85 26.41 9.08 9.86 0.55 6.82 6.58
Ours. Copy 22.67 23.54 37.51 37.45 29.76 19.01 28.32 8.80 10.51 0.58 6.68 6.64
Ours. Mix 23.21 25.33 38.71 40.64 31.33 19.35 29.76 9.46 10.66 0.61 6.92 6.91
@@
SKV 22.11 15.79 31.01 41.12 29.20 19.35 26.43 8.36 9.46 0.79 6.56 6.29
Ada-SKV 22.99 19.95 34.22 42.97 30.82 20.15 28.52 9.07 10.3 0.54 6.59 6.63
Ours. Copy 23.49 25.39 38.15 42.45 32.84 19.95 30.38 8.87 10.35 0.78 7.52 6.88
Ours. Mix 21.80 29.19 41.89 43.73 35.01 20.40 32.00 9.60 11.13 0.67 7.22 7.16
@@
SKV 23.38 20.18 37.65 42.80 33.23 20.01 29.54 9.04 10.59 0.53 7.53 6.92
PyramidKV 23.94 20.27 36.27 42.51 31.44 19.99 29.07 8.66 10.61 0.53 6.98 6.70
Ada-SKV 24.20 24.63 37.95 43.64 33.27 20.03 30.62 9.29 11.23 0.62 7.10 7.06
Ours. Copy 23.83 29.04 39.90 42.36 33.58 20.57 31.54 9.05 11.15 0.52 7.22 6.99
Ours. Mix 24.68 30.49 38.59 44.32 36.41 20.54 32.51 9.47 11.56 0.54 7.65 7.31
@@
SKV 25.47 23.75 38.64 43.66 33.98 19.83 30.89 9.00 11.07 0.63 7.34 7.01
PyramidKV 24.69 23.65 35.10 43.25 31.16 20.06 29.65 8.90 10.62 0.74 7.57 6.96
Ada-SKV 25.73 25.44 37.84 43.78 33.95 19.83 31.09 9.22 11.18 0.53 7.42 7.09
Ours. Copy 23.84 29.21 39.79 44.41 36.09 20.59 32.32 9.13 11.61 0.56 7.12 7.11
Ours. Mix 24.75 29.75 38.03 44.43 36.45 21.67 32.51 9.34 11.26 0.56 7.54 7.18
@@
SKV 25.76 27.50 38.38 43.40 34.81 20.07 31.65 9.61 11.34 0.53 7.22 7.18
Ada-SKV 25.79 29.24 38.74 43.93 36.34 19.79 32.31 8.65 11.41 0.53 7.71 7.08
Ours. Copy 24.85 30.94 39.82 43.52 36.58 20.37 32.68 9.20 11.67 0.55 7.71 7.28
Ours. Mix 24.66 30.82 39.56 43.97 36.47 22.24 32.95 9.02 11.51 0.47 7.85 7.21
@@
SKV 19.95 18.63 38.16 31.24 21.39 13.81 23.86 10.41 11.49 0.46 9.38 7.94
PyramidKV 20.91 19.61 38.05 32.18 22.87 15.26 24.81 10.64 11.69 0.56 9.06 7.99
Ada-SKV 22.70 21.28 42.39 34.35 22.40 14.85 26.33 10.56 11.50 0.45 8.72 7.81
Ours. Copy 24.23 25.22 46.02 38.82 26.05 17.41 29.63 10.94 13.14 0.63 9.11 8.46
Ours. Mix 21.77 26.57 48.39 40.12 26.76 16.21 29.97 11.19 13.94 0.48 9.87 8.87
@@
SKV 21.47 21.95 45.24 33.88 21.83 15.53 26.65 10.86 12.24 0.57 8.81 8.12
Ada-SKV 21.57 24.59 46.70 35.74 25.57 14.37 28.09 11.14 12.37 0.45 9.57 8.38
Ours. Copy 23.97 29.60 48.40 39.66 26.31 18.13 31.01 11.43 13.04 0.53 10.26 8.82
Ours. Mix 25.04 27.95 48.48 41.28 27.65 18.05 31.41 11.44 13.08 0.63 10.20 8.84
@@
SKV 22.26 24.94 48.30 36.76 25.16 14.93 28.72 11.07 12.39 0.53 9.13 8.28
PyramidKV 21.42 25.36 47.94 38.75 25.82 15.30 29.10 11.57 12.35 0.56 9.51 8.50
Ada-SKV 23.81 27.04 48.56 38.32 25.34 16.42 29.91 11.67 13.57 0.52 10.53 9.07
Ours. Copy 24.98 29.31 49.01 41.36 27.16 17.34 31.53 11.94 13.30 0.63 10.95 9.21
Ours. Mix 24.94 31.02 50.76 42.11 26.14 18.47 32.24 12.37 13.88 0.48 9.86 9.15
@@
SKV 24.18 28.87 48.74 38.84 25.48 15.04 30.19 11.96 13.47 0.52 10.50 9.11
PyramidKV 23.07 28.97 48.37 39.54 25.63 16.59 30.36 11.34 13.32 0.65 10.81 9.03
Ada-SKV 24.22 29.92 48.96 40.08 25.55 17.45 31.03 12.12 14.53 0.52 10.57 9.44
Ours. Copy 24.97 30.94 49.45 42.25 26.34 18.54 32.08 12.09 13.88 0.62 10.94 9.38
Ours. Mix 25.59 31.33 50.26 42.66 27.20 19.37 32.74 11.62 15.61 0.50 9.97 9.43
@@
SKV 25.38 30.22 49.29 41.84 26.60 18.08 31.90 11.69 13.89 0.52 10.54 9.16
Ada-SKV 24.82 31.49 48.80 41.18 27.38 18.22 31.98 11.96 13.82 0.53 9.92 9.06
Ours. Copy 25.87 31.44 49.55 41.95 27.09 19.88 32.63 12.21 14.17 0.50 10.58 9.37
Ours. Mix 25.64 32.54 50.49 41.80 27.88 18.89 32.87 11.94 14.93 0.50 10.49 9.47

### v4

@@ Abstract
Key-Value (KV) caching is a common technique to enhance the computational efficiency of Large Language Models (LLMs), but its memory overhead grows rapidly with input length. Prior work has shown that not all tokens are equally important for text generation, proposing layer-level KV cache compression to selectively retain key information. Recognizing the distinct roles of attention heads in generation, we propose HeadKV, a head-level KV cache compression method, and HeadKV-R2, which leverages a novel contextual reasoning ability estimation for compression. Our approach operates at the level of individual heads, estimating their importance for contextual QA tasks that require both retrieval and reasoning capabilities. Extensive experiments across diverse benchmarks (LongBench, LooGLE), model architectures (e.g., Llama-3-8B-Instruct, Mistral-7B-Instruct), and long-context abilities tests demonstrate that our head-level KV cache compression significantly outperforms strong baselines, particularly in low-resource settings (KV size = 64 & 128). Notably, our method retains just 1.5% of the KV cache while achieving 97% of the performance of the full KV cache on the contextual question answering benchmark. All code and data are available at https://github.com/FYYFU/HeadKV.
@@ 1 Introduction
We conduct experiments on various benchmarks requiring both retrieval and reasoning abilities, including QA tasks from LongBench (Bai et al., 2024) and LooGLE (Li et al., 2024b), using backbone models such as Llama-3-8B-Instruct (Dubey et al., 2024) and Mistral-7B-Instruct (Jiang et al., 2023a). Our results demonstrate that the head-level KV cache compression method outperforms previous approaches on nearly all tasks. Furthermore, the allocation strategy based on our estimated importance scores—reflecting both retrieval and reasoning abilities—outperforms allocation strategies based on retrieval ability alone. In challenging scenarios like needle-in-a-Haystack and reasoning-in-a-Haystack tests, our methods effectively preserve the model’s retrieval and reasoning capabilities. Finally, experiments on memory and latency reveal that our approach significantly reduces memory usage and decoding latency, comparing with the original full KV.
@@ 2.2 KV cache compression
Improving computational efficiency for LLMs, particularly for extremely long inputs, has attracted considerable research interest (Shazeer, 2019; Chen et al., 2023; Dao, 2024; Zhou et al., 2024; Ye et al., 2024), including caching key and value vectors in Multi-Head Attention to improve generation efficiency (Ge et al., 2024a; Xiao et al., 2024; Zhang et al., 2023; Li et al., 2024d; Cai et al., 2024). One challenge with full KV cache approaches is that the memory usage increases linearly with input length, leading to significant memory management challenges. Various KV cache compression techniques have been proposed to reduce memory usage and improve inference efficiency. For example, Xiao et al. (2024) addressed the ‘attention sink’ issue with StreamingLLM, retaining only the first $k$ tokens KV cache. Zhang et al. (2023) applied the Heavy Hitter Oracle strategy to select key cache entries, while Li et al. (2024d) used the attention scores of the last $\alpha$ tokens to identify relevant entries. Cai et al. (2024) introduced PyramidKV, assigning smaller cache budgets to higher layers based on attention matrix patterns.
@@ 3.1 Head-level Importance Score Estimation
[…]
where $N$ is the length of the inserted answer $\bm{k}$}, and ${\bm{a}}^{t}_{h}$ is the attention score on the combined input from head $h$ at $t$-sh decoding step. By using various settings of the inserted position $p_{i}$, they obtain the head-level Retrieval Heads distribution.
@@
where $a_{i}\in{\bm{a}}^{t}_{h}$ represents the $i$-th highest attention score from head $h$, and $\text{top-}i({\bm{a}}^{t}_{h})$ is the token with the $i$-th highest score at $t$-th decoding step. We compute the importance score by considering the entire correct answer and increasing the number of tokens evaluated per head (Importance Score Estimation in Figure 1). Intuitively, heads with higher attention on the correct answer $\bm{k}$} should receive higher importance scores. We further refine the score using attention weights, yielding a more accurate distribution, as shown in Eq. 2.
@@ 3.2 Head-Level KV Cache Allocation
Building on the head-level importance distributions, we propose a comprehensive KV cache allocation strategy. Each head $h$ is initially assigned a fixed KV cache size $b$ with an associated importance score $S_{h}$. To allow dynamic allocation, we create a shared budget pool $B$ by extracting a portion of the budget from each head, leaving the remainder as the basic budget. This process is illustrated in the Head-Level Allocation section of Figure 1. The budget pool $B$ is then distributed among the heads in proportion to their importance scores $S_{h}$. The importance score distribution $\bm{S}$} was L1-normalized to ensure that the sum of $S_{h}$ equals to 1. The final head-level KV cache allocation is as follows:
@@ 4.1 Experiment settings
We compare our head-level KV cache compression method against strong baselines using two open-source LLMs: Llama-3-8B-Instruct (Dubey et al., 2024) and Mistral-7B-Instruct (Jiang et al., 2023a). The evaluation is based on two benchmarks for long-context understanding: LongBench (Bai et al., 2024) and LooGLE (Li et al., 2024b). For LongBench, we use datasets from the Single-Doc QA and Multi-Doc QA categories to assess contextual reasoning. For LooGLE, we focus on the Long Dependency QA task, which includes four QA-related tasks. Dataset details are in Appendix Table 5.
@@ 4.4 Long-Context Retrieval and Reasoning
Figure 5: Needle-in-a-Haystack test results on Llama-3-8B-Instruct with KV cache = 128. We build our head-level KV cache method based on SnapKV and our proposed method significantly outperform all strong baselines. Moreover, our Retrieval-Reasoning Heads distribution maintains and improves long context retrieval ability. Results on Mistral-7B-Instruct can be found in Appendix Figure 8, which are consistent with results on Llama-3-8B-Instruct.
@@ 4.5 Memory & Latency
To assess the decoding latency of each method, we use 32K-length data from the Reasoning-in-a-Haystack test as input and set various generation lengths (1, 512, 1024, 2048, 4096) for comparison. As shown in the Decoding Latency of Figure 6, our proposed method achieves the same decoding latency as other KV cache compression methods while maintaining performance closest to the Full KV cache. Notably, the decoding latency includes both the pre-filling time and the decoding time. Therefore, we can conclude that the pre-filling time for our method and other baselines (such as PyramidKV, Ada-SnapKV) is almost negligible, as shown at the starting point (generation length = 1) of Decoding Latency in Figure 6.
@@ Appendix A Detaile Results
Method Single-Doc QA Multi-Doc QA Avg. Long dependency QA Avg. $\beta$
@@
Llama-3-8B-Instruct, KV Size=Full -
FKV 25.56 32.07 39.71 43.57 35.28 21.18 32.90 8.73 11.21 0.67 7.43 7.01 -
@@
SKV 20.51 12.80 31.69 37.02 25.91 17.02 24.16 8.84 9.43 0.66 6.18 6.28 -
PyramidKV 21.17 13.66 29.34 34.86 23.46 15.88 23.06 8.27 9.31 0.63 6.86 6.27 -
Ada-SKV 22.26 17.30 33.37 39.82 27.86 17.85 26.41 9.08 9.86 0.55 6.82 6.58 -
Ours. Copy 22.67 23.54 37.51 37.45 29.76 19.01 28.32 8.80 10.51 0.58 6.68 6.64 2
Ours. Mix 23.21 25.33 38.71 40.64 31.33 19.35 29.76 9.46 10.66 0.61 6.92 6.91 1.2
@@
SKV 22.11 15.79 31.01 41.12 29.20 19.35 26.43 8.36 9.46 0.79 6.56 6.29 -
PyramidKV 22.01 17.05 31.52 39.27 28.99 18.34 26.20 8.89 9.63 0.61 6.72 6.46 -
Ada-SKV 22.99 19.95 34.22 42.97 30.82 20.15 28.52 9.07 10.3 0.54 6.59 6.63 -
Ours. Copy 23.49 25.39 38.15 42.45 32.84 19.95 30.38 8.87 10.35 0.78 7.52 6.88 1.5
Ours. Mix 21.80 29.19 41.89 43.73 35.01 20.40 32.00 9.60 11.13 0.67 7.22 7.16 1.01
@@
SKV 23.38 20.18 37.65 42.80 33.23 20.01 29.54 9.04 10.59 0.53 7.53 6.92 -
PyramidKV 23.94 20.27 36.27 42.51 31.44 19.99 29.07 8.66 10.61 0.53 6.98 6.70 -
Ada-SKV 24.20 24.63 37.95 43.64 33.27 20.03 30.62 9.29 11.23 0.62 7.10 7.06 -
Ours. Copy 23.83 29.04 39.90 42.36 33.58 20.57 31.54 9.05 11.15 0.52 7.22 6.99 1.1
Ours. Mix 24.68 30.49 38.59 44.32 36.41 20.54 32.51 9.47 11.56 0.54 7.65 7.31 1.1
@@
SKV 25.47 23.75 38.64 43.66 33.98 19.83 30.89 9.00 11.07 0.63 7.34 7.01 -
PyramidKV 24.69 23.65 35.10 43.25 31.16 20.06 29.65 8.90 10.62 0.74 7.57 6.96 -
Ada-SKV 25.73 25.44 37.84 43.78 33.95 19.83 31.09 9.22 11.18 0.53 7.42 7.09 -
Ours. Copy 23.84 29.21 39.79 44.41 36.09 20.59 32.32 9.13 11.61 0.56 7.12 7.11 1.2
Ours. Mix 24.75 29.75 38.03 44.43 36.45 21.67 32.51 9.34 11.26 0.56 7.54 7.18 1.1
@@
SKV 25.76 27.50 38.38 43.40 34.81 20.07 31.65 9.61 11.34 0.53 7.22 7.18 -
PyramidKV 25.38 26.83 36.90 44.09 34.24 21.49 31.49 8.98 11.41 0.53 6.96 6.97 -
Ada-SKV 25.79 29.24 38.74 43.93 36.34 19.79 32.31 8.65 11.41 0.53 7.71 7.08 -
Ours. Copy 24.85 30.94 39.82 43.52 36.58 20.37 32.68 9.20 11.67 0.55 7.71 7.28 1.2
Ours. Mix 24.66 30.82 39.56 43.97 36.47 22.24 32.95 9.02 11.51 0.47 7.85 7.21 1.2
@@
SKV 19.95 18.63 38.16 31.24 21.39 13.81 23.86 10.41 11.49 0.46 9.38 7.94 -
PyramidKV 20.91 19.61 38.05 32.18 22.87 15.26 24.81 10.64 11.69 0.56 9.06 7.99 -
Ada-SKV 22.70 21.28 42.39 34.35 22.40 14.85 26.33 10.56 11.50 0.45 8.72 7.81 -
Ours. Copy 24.23 25.22 46.02 38.82 26.05 17.41 29.63 10.94 13.14 0.63 9.11 8.46 1.5
Ours. Mix 21.77 26.57 48.39 40.12 26.76 16.21 29.97 11.19 13.94 0.48 9.87 8.87 1.2
@@
SKV 21.47 21.95 45.24 33.88 21.83 15.53 26.65 10.86 12.24 0.57 8.81 8.12 -
PyramidKV 21.76 21.98 43.72 32.76 22.73 15.59 26.42 10.64 11.9 0.47 8.69 7.93 -
Ada-SKV 21.57 24.59 46.70 35.74 25.57 14.37 28.09 11.14 12.37 0.45 9.57 8.38 -
Ours. Copy 23.97 29.60 48.40 39.66 26.31 18.13 31.01 11.43 13.04 0.53 10.26 8.82 1.5
Ours. Mix 25.04 27.95 48.48 41.28 27.65 18.05 31.41 11.44 13.08 0.63 10.20 8.84 1.1
@@
SKV 22.26 24.94 48.30 36.76 25.16 14.93 28.72 11.07 12.39 0.53 9.13 8.28 -
PyramidKV 21.42 25.36 47.94 38.75 25.82 15.30 29.10 11.57 12.35 0.56 9.51 8.50 -
Ada-SKV 23.81 27.04 48.56 38.32 25.34 16.42 29.91 11.67 13.57 0.52 10.53 9.07 -
Ours. Copy 24.98 29.31 49.01 41.36 27.16 17.34 31.53 11.94 13.30 0.63 10.95 9.21 2
Ours. Mix 24.94 31.02 50.76 42.11 26.14 18.47 32.24 12.37 13.88 0.48 9.86 9.15 1.005
@@
SKV 24.18 28.87 48.74 38.84 25.48 15.04 30.19 11.96 13.47 0.52 10.50 9.11 -
PyramidKV 23.07 28.97 48.37 39.54 25.63 16.59 30.36 11.34 13.32 0.65 10.81 9.03 -
Ada-SKV 24.22 29.92 48.96 40.08 25.55 17.45 31.03 12.12 14.53 0.52 10.57 9.44 -
Ours. Copy 24.97 30.94 49.45 42.25 26.34 18.54 32.08 12.09 13.88 0.62 10.94 9.38 1.01
Ours. Mix 25.59 31.33 50.26 42.66 27.20 19.37 32.74 11.62 15.61 0.50 9.97 9.43 1.005
@@
SKV 25.38 30.22 49.29 41.84 26.60 18.08 31.90 11.69 13.89 0.52 10.54 9.16 -
PyramidKV 24.28 30.05 49.17 40.49 26.43 18.80 31.54 11.77 14.51 0.51 10.19 9.25 -
Ada-SKV 24.82 31.49 48.80 41.18 27.38 18.22 31.98 11.96 13.82 0.53 9.92 9.06 -
Ours. Copy 25.87 31.44 49.55 41.95 27.09 19.88 32.63 12.21 14.17 0.50 10.58 9.37 1.5
Ours. Mix 25.64 32.54 50.49 41.80 27.88 18.89 32.87 11.94 14.93 0.50 10.49 9.47 1.01